OpenStudy (anonymous):

check the answer plse: divide and simplify 9/x divided by 36/x = 4/x^2 answer

OpenStudy (anonymous):

Incorrect: \[9/x \div 36/x = 9/x * x/36 = 9x/36x = 1/4\]